OpenStudy (anonymous):

What is the equation of the line through the point (2,4) with a slope of 3? a. y = 3x + 4 b. y = 3x + 10 c. y = 3x - 2 d. y = 2x + 4

OpenStudy (aroub):

y=mx+b (2,4) 4=3(2)+b 4=6+b -6+4=b -2=b so the answer is y=3x-2 (c.)

OpenStudy (anonymous):

answer c, 3x-y-2=0